# A meaningful prediction of functional decline in amyotrophic lateral sclerosis based on multi-event survival analysis

### v1

@@ Introduction
Amyotrophic lateral sclerosis (ALS) is a neurological disease that causes a gradual loss of upper and lower motor neurons in the central nervous system. Unfortunately, most patients will die within 2-5 years after onset [1], primarily due to complications from respiratory failure [2]. The underlying causes of the disease are not yet understood, and to this day, current treatments aim to prolong survival and improve quality of life. Moreover, due to its heterogeneous and unpredictable nature, and the high variability in its progression rate and clinical phenotype [5], it is often difficult to determine the appropriate treatment. This also makes it difficult to determine the optimal time for such medical interventions, e.g., noninvasive ventilation, or if a specific treatment can effectively slow the disease progression. Therefore, there is a significant need for tools that can predict disease progression, facilitate treatment plans, and improve patient care.
@@
The ALS Functional Rating Scale-Revised (ALSFRS-R) [6] is the most widely-used questionnaire to evaluate the progression of ALS (see Table 1). It contains 12 questions to assess functional capacity in patients, with a focus on bulbar, motor, and respiratory functions – e.g., swallowing, walking, and breathing, respectively. Each response is scored between 0 and 4, where 0 represents complete inability with regard to that function, and 4 represents normal function.Based on this questionnaire, an interesting task is to predict the degree of functional decline a patient experiences during the course of the disease, rather than simply when the patient is likely to die [7, 8]. Many previous projects have attempted to predict the future ALSFRS-R score based on patient features (covariates) [12, 9, 10, 11, 3]. Ong et al. [12] proposed a binary model to predict the functional decline class (i.e., whether ALS progresses slowly or rapidly in a specific patient) after baseline (trial entry), based on several patient characteristics. Their model was able to predict slow or rapid decline, but did not predict when functional decline would occur over time. Similarly, Jabbar et al. [11] proposed a method to predict disease progression as fast or slow, but as in [12], their model could only give a yes/no answer to whether a patient would experience functional decline over time. Amaral et al. [10] applied several clustering algorithms to clinical patient records, which revealed four separate groups regarding disease progression. Their method could assign newly-diagnosed patients to some progression group, but did not predict when a patient would experience functional decline based on their covariates. Finally, Vieira et al. [9] trained two machine learning models to predict a patient’s future ALSFRS-R score: a voice model based on speech recordings and an accelerometer model based on accelerometer measurements. These models showed promise in predictive performance, but similar to [10] and others, they simply predicted what the ALSFRS-R score would be for a patient in the future, not the time until functional decline would occur. In addition, disease progression is not necessarily constant or linear in ALS, which the authors assumed, but can vary significantly over time and between individuals [4].
@@
Based on these shortcomings in the literature, we propose a novel method that can predict when an ALS patient will experience each of five common types of functional decline, based on the ALSFRS-R protocol (see Table 1): significant difficulty ($\leq$ 2) in each of Speaking, Swallowing, Handwriting, Walking and Dyspnea. The events are distinct but not mutually exclusive and likely share some information in their covariates; it may be that Mr. Smith with limb-onset ALS will experience difficulty walking and writing earlier than Mr. Johnson, who has bulbar-onset ALS. On the other hand, Mr. Johnson may have difficulty speaking and swallowing solid foods before Mr. Smith. As a newly-diagnosed patient with ALS, our approach will help answer important questions, such as:
@@
In practice, we formulate the problem as a multi-event survival problem: given an instance (e.g., a description of Mr. Smith at a “baseline” point in time), our method predicts the time until said Mr. Smith will experience each of the events listed above. We present each prediction as individual survival (event) distributions (ISDs) [13], which provide the probability of each event at each future time point, after the baseline visit. These ISDs were recently used to predict mortality in patients with ALS [7]. We describe each patient by the values of a set of covariates such as age, site of onset, forced vital capacity (FVC), and others, which are recorded in clinical records obtained at or before the patient’s current visit (baseline). To our knowledge, no existing method can predict the time of distinct individual events in ALS that explicitly provide the time. Fig. 1 shows an outline of the proposed method. The source code is publicly available at: https://github.com/thecml/FunctionalALS.
{adjustwidth}
-2.25in0in Refer to caption
@@
Figure 1: Outline of the proposed method. Historical data about ALS patients form a patient dataset $\mathcal{D}$ with $N$ training instances and $d$ covariates, and the time to event (filled dot) or censoring (hollow dot) from a patient’s first visit – the “baseline time”. We consider five separate, but related events, i.e.Speech, Swallowing, Handwriting, Walking, and Dyspnea. We then use the recorded covariates (taken at the baseline time) and event information to train a survival model $\mathcal{M}$ that can accurately estimate the individual survival distribution (ISD) of each of these five events, for a novel patient $\bm{x}_{i}$, denoted as $\hat{S}^{(i)}$. These ISDs give the probability of each of these five events occurring after $t$ days after the baseline visit, for all $t>0$. These ISDs can then be used to estimate the time to event for this $\bm{x}_{i}$, for example when the survival (event) curve intersects the dashed horizontal line at 50%, for each of the five events.
{adjustwidth}
-2.25in0in ID Domain Assessment Description Score 1 Bulbar Speech Ability to speak clearly and be understood. 4 = Speaks clearly; 2 = Intelligible with repeating; 0 = Loss of useful speech 2 Bulbar Salivation Control over saliva and drooling. 4 = Normal salivation; 2 = Moderately excessive saliva; 0 = Marked drooling 3 Bulbar Swallowing Ability to swallow without choking. 4 = Normal eating habits; 2 = Difficulty with some foods; 0 = Needs feeding tube 4 Fine motor Handwriting Ability to write or type. 4 = Normal writing; 2 = Not all words legible; 0 = Unable to grip pen 5 Fine motor Cutting food and handling utensils Ability to cut food and use utensils independently. 4 = Normal; 2 = Can cut most foods, but slow; 0 = Needs to be fed 6 Fine motor Dressing and hygiene Ability to dress and complete personal hygiene. 4 = Fully independent; 2 = Requires some assistance; 0 = Fully dependent 7 Gross motor Turning in bed Ability to move in bed. 4 = Normal movement; 2 = Can turn alone, but slow; 0 = Fully dependent 8 Gross motor Walking Ability to walk without assistance. 4 = Walks normally; 2 = Walks with assistance; 0 = Cannot walk 9 Gross motor Climbing stairs Ability to climb stairs. 4 = Climbs normally; 2 = Mild unsteadiness or fatigue; 0 = Cannot climb stairs 10 Respiratory Dyspnea Shortness of breath. 4 = None; 2 = Moderate shortness of breath; 0 = Severe shortness of breath 11 Respiratory Orthopnea Ability to breathe easily when lying down. 4 = None; 2 = Requires elevated pillows; 0 = Requires ventilatory support 12 Respiratory Respiratory insufficiency Need for ventilatory support. 4 = None; 2 = Requires non-invasive ventilation; 0 = Requires invasive ventilation
@@
Table 1: The definition of the 12 functional assessments in the ALS Function-Rating-Scale Revised (ALSFRS-R) test [6] and their scores. Each score is from 0 to 4 points, where 0 represents complete inability with regard to the function, and 4 represents normal function. Bold highlights the five assessments used in this work. For each patient, our method predicts the time until that patient will experience significant difficulty (ALSFRS-R $\leq$ 2) as regards to the specific assessment.
@@
For our empirical analyses, we use the publicly-available Pooled Resources Open-Access Clinical Trials (PRO-ACT) [^1] Data used in the preparation of this article were obtained from the Pooled Resource Open-Access ALS Clinical Trials (PRO-ACT) Database. The data available in the PRO-ACT Database have been volunteered by PRO-ACT Consortium members. The dataset was obtained on October 1st, 2024. [14] dataset, which is the largest ALS dataset in the world, including patient demographics, lab and medical records and family history of over 11,600 ALS patients from 23 clinical trials (see Table 2). In each clinical trial, patients follow a schedule of visits until the end of the study or until they become too ill to participate. Covariates are recorded at each visit, which include the total ALSFRS-R score, age, sex, region of onset (e.g., limb, bulbar), the mean FVC score, time since diagnosis, disease progression rate [^2] Defined as (48 - the total ALSFRS-R score)/(days since diagnosis/30), and whether or not the patient is taking Riluzole. For our analysis, we only use covariates recorded at the baseline visit for each individual patient.
To define our events of interest, each patient’s disease state is assessed by a clinician at every visit using the ALSFRS-R scale. In this study, we focus on five specific events: Speech, Swallowing, Handwriting, Walking, and Dyspnea. To annotate our dataset, we regard an event to have occurred the first time a patient scores 2 or less in that specific assessment during any follow-up visit following their baseline visit. These five events can occur in any order and are not mutually exclusive. There are several types of censoring, including death, being unable to perform the assessment or leaving the study prematurely. We use a maximum follow-up time of 500 days to focus on the early stages of the disease and because data becomes sparse with high censoring afterward. We consider patients who score 2 or less on any of the five assessments at baseline as having already had the event, indicating that the loss of function occurred before the start of the trial, and exclude these patients from the study. We also exclude patients with no recorded ALSFRS-R history. This makes a dataset of 3,220 patients. The code for loading and annotating the data is available in the source code repository. Table 3 is an overview of the dataset. Fig. 2 shows the distribution of event and censoring times. A list of covariates is available in the Supplement (S1 Table.).
@@
Demographic/Data PRO-ACT (N=3,220) Pct. or mean ($\pm$ SD)
@@
Table 3: Overview of the dataset. “SP”, “SW” “HA”, “WA” and “DY”, are the Speech, Swallowing, Handwriting, Walking and Dyspnea events, respectively. Event distribution indicates the percentage of uncensored instances of each type in the dataset. $N$: sample size, $d$: number of covariates, $K$: number of event types, $T$: observed time (of the uncensored instances) for any event in days.
Dataset N d K Tmin Tmax Tmean Event distribution (%):
PRO-ACT 3,220 8 5 1 498 130 SP: 37.8, SW: 31.9 HA: 50.3, WA: 60.6 DY: 27.0
{adjustwidth}
-2.25in0in Refer to caption
@@
Survival analysis models the time until some event of interest occurs. This event can either be observed or censored – the latter typically happens if the patient is still alive at the end of the study or decides to drop out [15, Ch. 11]. Survival analysis has become an important tool for understanding diseases and predicting outcomes, helping medical researchers evaluate the significance of prognostic factors in applications like comatose after cardiac arrest [16], liver transplantation [17] or fall risk [18]. In ALS, survival analysis has been used mainly to predict risk scores with respect to death [4, 7, 19].
We briefly introduce the notation: Let $\mathcal{D}=\left\{\left(\bm{x}^{(i)},\bm{t}^{(i)},\bm{\delta}^{(i)}\right)\right\}_{i=1}^{N}$ be the dataset, where for each $i$, $\bm{x}^{(i)}\in\mathbb{R}^{d}$ is a $d$-dimensional vector of covariates, and we consider discrete times: $\bm{t}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}^{K}$ is a vector of observed times for the $i$-th patient for $K$ different events, each at one of these $T_{\text{max}}$ possible times, and $\bm{\delta}^{(i)}\in\{0,1\}^{K}$ is the vector of event indicators for each event. Moreover, let $e_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the event time and let $c_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the censoring time for the $i$-th instance for event $k$ over the event horizon, thus $t_{k}^{(i)}=e_{k}^{(i)}$ if $\delta_{k}^{(i)}=1$, otherwise, $t_{k}^{(i)}=c_{k}^{(i)}$ if $\delta_{k}^{(i)}=0$. A survival model predicts the probability that some event $k$ occurs at time $T$ later than $t$, i.e., the survival probability, $S\left(t\right)=\text{Pr}\left(T>t\right)=1-\text{Pr}\left(t\leq T\right)$. In multi-event survival analysis, we want to predict the probability of different events occurring over time. For each patient $i$, this results in a matrix of survival functions, $\bm{S}^{(i)}=\{\bm{s}_{1}^{(i)},\bm{s}_{2}^{(i)},\ldots,\bm{s}_{K}^{(i)}\}$. Here, $\bm{s}_{k}^{(i)}$ represents the probabilities that patient $i$ does not experience event $k$ up to each time point $t$, i.e.$\bm{s}_{k}^{(i)}=(P(e_{k}^{(i)}>1),P(e_{k}^{(i)}>2),\ldots,P(e_{k}^{(i)}>T_{\text{max}}))$. See Fig. 1 for a visual example.
@@
We follow Sechidis et al. [20] and first split the PRO-ACT dataset into a training, validation and test set by 70%, 10%, and 20% using a stratified procedure, balanced to ensure that event times and censoring rates are consistent across the three datasets and the $K$ events. The stratification procedure uses a random seed (0-9). After splitting the data, we impute missing values using sample mean for real-valued covariates or mode for categorical covariates based on the training set only. A list of missing values is available in the Supplement (S2 Table.). After imputing missing values, we encode categorical covariates using a one-hot encoding strategy and apply a $z$-score data normalization to speed up the training process and improve stability. We do not perform any outlier detection.
Our evaluation procedure consists of training five survival models on the training set and evaluating them on the holdout test set. For evaluation fairness, we do not perform hyperparameter optimization for specific models, but configure all models with sensible default parameters. A list of hyperparameters is available in the Supplement (S3 Table.). For deep learning models, we use a single hidden layer with 32 nodes. For methods that support it, we use early stopping during training to terminate the process if the validation loss does not improve for 10 consecutive epochs. All data preprocessing and experiments were implemented in Python 3.9. A list of software packages used is available in the source code repository.
@@
For the purpose of comparability in the results, we implement five survival models to validate our method. As single-event models, we implement the traditional Cox Proportional Hazards (CoxPH) model using a linear estimator [21], Random Survival Forests (RSF) [22], DeepSurv [23] and MTLR [24]. As a multi-event model, we implement the Multi-Event Network for Survival Analysis (MENSA) model [25]. To explain these models: CoxPH is a popular semiparametric method to fit a regression model to survival data. By adopting a multiplicative form for the contribution of several covariates to each individual’s event time, the CoxPH model is a powerful yet simple tool for assessing the simultaneous effect different covariates have on event times. RSF extends decision trees for survival analysis and recursively splits training data based on some survival-specific criterion, e.g., logrank splitting, with the goal of maximizing separation between event times between nodes. RSF can model nonlinear relationships between the covariates and the event and does not assume proportional hazards. DeepSurv is a multilayer perceptron (MLP) based on the CoxPH model, where the risk score is a nonlinear function of the covariates. DeepSurv assume proportional hazards. MTLR is a discrete survival model that estimates the individual survival function as a multinomial logistic regression model and does not assume proportional hazards. Lastly, MENSA is an MLP that models the individual survival distribution for $K$ events as a mixture of Weibull distributions with a shared covariate layer to model dependencies between the events. MENSA does not assume proportional hazards.
@@
To comprehensively assess the survival models, we select a range of evaluation metrics that capture three key aspects of model performance: discrimination, absolute/squared error, and calibration. Discrimination, usually measured by the concordance index (CI), is relevant when we want to identify which patients are likely to experience functional decline earlier versus later at the group level. Squared or absolute error metrics, like the Brier Score (BS) and mean absolute error (MAE), are used to assess how accurately the model predicts the time of functional impairment for individuals. Finally, calibration measures how well the predicted event probabilities match the observed frequencies of events at specific time points. For example, if a model predicts that a patient has a 80% chance of not experiencing significant functional decline before 100 days, then among 100 similar patients, about 80 should not experience significant functional decline by that time. We use the following metrics to assess discrimination, absolute/squared error, and calibration:
@@
CI: Harrell’s concordance index [26] measures discrimination performance by calculating the proportion of concordant pairs among all comparable pairs given predicted risk scores. A pair is considered comparable if the event order can be determined. Specifically, for a comparable pair $\{i,j\}$ with event times $y_{i}<y_{j}$ and $\delta_{i}=1$ (indicating that $i$ experienced the event), we calculate risk scores $\hat{\eta}_{i}$ and $\hat{\eta}_{j}$ as the negative median survival times. If the predicted risk score for $j$ is greater than that for $i$ at the time $i$ experienced the event (while $j$ remains event-free), the pair is concordant. A concordance index of 0.5 indicates chance-level ranking. We also report Uno’s CI [27], which is a concordance index that uses a technique called “inverse probability of censoring weighting”, or simply IPCW, to correct for the bias introduced by censoring.
@@
BS/IBS: The Brier Score calculates the squared error between the predicted probability of the event and the Heaviside step function of the observed event. The Integrated Brier Score [28] (IBS) aggregates Brier Scores across multiple time points to provide a single measure of the squared error. We use IPCW to weigh censored instances accordingly.
@@
MAE: The mean absolute error calculates the average absolute difference between the true event time and the predicted event time. Given an individual survival distribution, $S\left(t\mid\bm{x}_{i}\right)=\text{Pr}\left(T>t\mid\bm{x}_{i}\right)$, we calculate the predicted event time $\hat{t}_{i}$ as the median survival time [29]. For censored individuals, we calculate the marginal MAE (mMAE) as proposed by [13], which assigns a “best-guess” estimate to each censored individual using the Kaplan-Meier (KM) [30] estimator.
@@
D-calibration: Distribution calibration [13], aka D-Cal, measures how well the predicted survival function $\hat{S}(t)$ is calibrated for each event. We assess this using Pearson’s chi-squared test with $\alpha=0.05$. For any probability interval $[a,b]\in[0,1]$, we define $D_{m}(a,b)$ as the group of individuals in the dataset $D$ whose predicted probability of an event is in the interval $[a,b]$ [31]. A model is D-calibrated if the proportion of individuals $|D_{m}(a,b)|/|D|$ is statistically similar to the amount $b-a$. In our results, we report the number of times each model is D-calibrated (at $p>0.05$) for each event across ten experiments.
@@
We present our empirical results (Table 4 and Fig. 3) in terms of three key aspects with respect to performance: discrimination, absolute/squared error, and calibration. Below, we discuss the results and critically reflect on their implications.

### v2

@@ Introduction
Amyotrophic lateral sclerosis (ALS) is a neurological disease that causes a gradual loss of upper and lower motor neurons in the central nervous system. Unfortunately, most patients will die within 2-5 years after onset [1], primarily due to complications from respiratory failure [2]. The underlying causes of the disease are not yet understood, and as there are no known cures, current treatments aim only to prolong survival and improve quality of life. Moreover, due to its heterogeneous and unpredictable nature, and the high variability in its progression rate and clinical phenotype [5], it is often difficult to determine the appropriate treatment. This also complicates decisions about the optimal timing of medical interventions, e.g., non-invasive ventilation, or if a specific treatment can effectively slow the progression of the disease. There is therefore a significant need for tools that can predict disease progression to facilitate personalized treatment plans and improve patient care.
@@
The ALS Functional Rating Scale-Revised (ALSFRS-R) [6] is the most widely used questionnaire to evaluate the progression of ALS. It contains 12 questions to assess functional capacity in patients, with a focus on bulbar, motor, and respiratory functions – e.g., swallowing, walking, and breathing, respectively. Each response is scored between 0 and 4, where 0 represents complete inability with regard to that function, and 4 represents normal function. As this questionnaire is administered to patients at multiple time points over the course of the disease, it provides a longitudinal view of functional decline across several areas. Based on this questionnaire, an interesting task is to predict the degree of functional decline a patient experiences during the course of the disease, rather than simply predicting when a patient is likely going to die [7, 8]. Many previous projects have therefore attempted to predict the future ALSFRS-R score based on patient covariates [12, 9, 10, 11, 3]. Ong et al. [12] proposed a binary model to predict the functional decline class (i.e., whether ALS progresses slowly or rapidly in a specific patient) after baseline (trial entry), based on several patient characteristics. Their model was able to predict slow or rapid decline, but did not predict when functional decline would occur over time. Similarly, Jabbar et al. [11] proposed a method to predict disease progression as fast or slow, but as in [12], their model could only give a yes/no answer to whether a patient would experience functional decline over time. Amaral et al. [10] applied several clustering algorithms to clinical patient records, which revealed four separate groups regarding disease progression. Their method could assign newly-diagnosed patients to a progression group, but did not predict when a patient would experience functional decline based on their covariates. Finally, Vieira et al. [9] trained two machine learning models to predict a patient’s future ALSFRS-R score: a voice model based on speech recordings and a movement model based on accelerometer measurements. These models showed strong predictive performance, but similar to Amaral et al. [10] and others, they predicted future ALSFRS-R scores rather than estimating the time until functional decline would occur. Moreover, disease progression in ALS is neither constant nor linear, as those studies implicitly assumed, but can vary substantially over time and between individuals [4].
@@
Based on these shortcomings, we propose a novel method that can predict when an ALS patient will experience each of five common types of functional decline, based on the ALSFRS-R protocol (see Table 1): significant difficulty ($\leq$ 2) in each of Speaking, Swallowing, Handwriting, Walking, and Dyspnea. These five events are distinct but not mutually exclusive and likely share some information in their covariates; it may be that a patient (e.g., Mr. Smith), with limb-onset ALS, will experience difficulty walking and writing earlier than another patient (e.g., Mr. Johnson), who has bulbar-onset ALS. On the other hand, Mr. Johnson may have difficulty speaking and swallowing solid foods before Mr. Smith. As a newly-diagnosed patient with ALS, our approach will help answer important questions, such as:
@@
In practice, we formulate the problem as a multi-event survival problem: given an instance (e.g., a description of Mr. Smith at a “baseline” point in time), our method predicts the time until said Mr. Smith will experience each of the events listed above. We present each prediction as an individual survival distribution (ISD) [13], which provide the probability of each event at each future time point, after the baseline visit. We describe each patient by the values of a set of covariates such as age, site of onset, forced vital capacity (FVC), and others, which are recorded in clinical records obtained at or before the baseline visit. To our knowledge, no existing method can explicitly predict the timing of distinct individual events in ALS. Fig. 1 shows an outline of the proposed method. The source code is publicly available at: https://github.com/thecml/FunctionalALS.
[Fig. 1]
@@
Figure 1: Outline of the proposed method. Historical data about ALS patients form a patient dataset $\mathcal{D}$ with NN training instances and $d$ covariates, and the time to event (filled dot) or censoring (hollow dot) from a patient’s first visit – the “baseline time”. Censoring indicates that the event of interest was not observed for a patient within the study period, so their exact event time remains unknown. We consider five separate but related events, i.e., Speech, Swallowing, Handwriting, Walking, and Dyspnea. We use the recorded covariates (taken at the baseline time) and event information to train a survival model $\mathcal{M}$ that can accurately estimate the individual survival distribution (ISD) of each of these five events, for a novel patient $\bm{x}_{i}$, denoted as $\hat{S}^{(i)}$. These ISDs give the probability of each of these five events occurring after $t$ days after the baseline visit, for all $t>0$. They can also be used to estimate the time to event for this $\bm{x}_{i}$ patient, for example, when the survival curve intersects the dashed horizontal line at 50%, which is called the median survival time, for each of the five events.
@@
Table 1: The definition of the five included functional assessments from the ALS Functional Rating Scale-Revised (ALSFRS-R) [6] and their scores. Each score ranges from 0 to 4, where 0 represents complete inability with regard to the function, and 4 represents normal function. For each patient, our method predicts the time until that patient will experience significant difficulty (ALSFRS-R $\leq$ 2) for that specific assessment.
Domain
Assessment
Description
Score
Bulbar
Speech
Ability to speak clearly and be understood.
4 = Speaks clearly;
2 = Intelligible with repeating;
0 = Loss of useful speech
Bulbar
Swallowing
Ability to swallow without choking.
4 = Normal eating habits;
2 = Difficulty with some foods;
0 = Needs feeding tube
Fine motor
Handwriting
Ability to write or type.
4 = Normal writing;
2 = Not all words legible;
0 = Unable to grip pen
Gross motor
Walking
Ability to walk without assistance.
4 = Walks normally;
2 = Walks with assistance;
0 = Cannot walk
Respiratory
Dyspnea
Shortness of breath.
4 = None;
2 = Moderate shortness of breath;
0 = Severe shortness of breath
@@
For our empirical analyses, we use the publicly-available Pooled Resources Open-Access Clinical Trials (PRO-ACT) [^1] Data used in the preparation of this article were obtained from the Pooled Resource Open-Access ALS Clinical Trials (PRO-ACT) Database. The data available in the PRO-ACT Database have been volunteered by PRO-ACT Consortium members. The dataset was obtained on October 1st, 2024. [14] dataset, which is the largest ALS dataset in the world. It includes patient demographics, lab and medical records, as well as family histories, of over 11,600 ALS patients from 23 clinical trials (see Table 2). In each clinical trial, patients follow a schedule of visits until the end of the study or until they become too ill to participate. Covariates recorded at each visit include the total ALSFRS-R score, age, sex, region of onset (e.g., limb, bulbar), the mean FVC score, time since diagnosis, disease progression rate [^2] Defined as (48 - the total ALSFRS-R score)/(days since diagnosis/30)., and whether or not the patient is taking Riluzole. In this study, we only use covariates recorded at the baseline visit for each individual patient.
To define our events of interest, each patient’s disease state is assessed by a clinician at every visit using the ALSFRS-R scale. We focus on five specific events: Speech, Swallowing, Handwriting, Walking, and Dyspnea. To annotate our dataset, we define an event as having occurred the first time a patient scores 2 or fewer in that specific assessment during any follow-up visit following their baseline visit. These five events can occur in any order and are not mutually exclusive. There are several types of censoring, including death, being unable to perform the assessment or leaving the study prematurely. Censoring thus represents incomplete event information, where the true event time is only known to exceed the observed follow-up time. We use a maximum follow-up time of 500 days to focus on the early stages of the disease and because data become sparse with high censoring afterward. We consider patients who score 2 or fewer on any of the five assessments at baseline as having already had the event, indicating that the loss of function occurred before the start of the trial, and exclude these patients from the study. We also exclude patients with no recorded ALSFRS-R history. After applying these criteria and accounting for censoring, the final dataset comprises 3,220 patients with observed or right-censored event times. The code for loading and annotating the data is available in the source code repository. Table 3 provides an overview of the dataset. Fig. 2 shows the distribution of event and censoring times. A list of covariates is available in the Supplement (S1 Table.).
@@
Demographic/Data
PRO-ACT (N=3,220)
Pct. or mean ($\pm$ SD)
@@
Table 3: Overview of the dataset. “SP”, “SW”, “HA”, “WA” and “DY”, are the Speech, Swallowing, Handwriting, Walking and Dyspnea events, respectively. Event distribution indicates the percentage of uncensored instances of each type in the dataset. NN: sample size, $d$: number of covariates, $K$: number of event types, $T$: observed time (of the uncensored instances) for any event in days.
Dataset N d
K
Tmin
Tmax
Tmean
Event distribution (%):
PRO-ACT 3,220 8 5 1 498 130
SP: 37.8, SW: 31.9
HA: 50.3, WA: 60.6
DY: 27.0
[Fig. 2]
@@
Survival analysis models the time until some event of interest occurs. This event can either be observed or censored – the latter typically happens if the patient is still alive at the end of the study or decides to drop out [15, Ch. 11]. Survival analysis has become an important tool for understanding diseases and predicting outcomes, which helps medical researchers evaluate the significance of prognostic factors in applications like comatose after cardiac arrest [16], liver transplantation [17] or fall risk [18]. In ALS, survival analysis has been used mainly to predict risk scores with respect to death [4, 7, 19].
We briefly introduce the notation: Let $\mathcal{D}=\left\{\left(\bm{x}^{(i)},\bm{t}^{(i)},\bm{\delta}^{(i)}\right)\right\}_{i=1}^{N}$ be the dataset, where for each $i$, $\bm{x}^{(i)}\in\mathbb{R}^{d}$ is a $d$-dimensional vector of covariates. We consider discrete times, where $\bm{t}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}^{K}$ is a vector of observed times for the $i$-th patient for $K$ different events, each at one of these $T_{\text{max}}$ possible times, and $\bm{\delta}^{(i)}\in\{0,1\}^{K}$ is a vector of event indicators for each event. Moreover, let $e_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the event time and let $c_{k}^{(i)}\in\{1,2,\ldots,T_{\text{max}}\}$ denote the censoring time for the $i$-th instance for event $k$ over the event horizon, thus we have $t_{k}^{(i)}=e_{k}^{(i)}$ if $\delta_{k}^{(i)}=1$, otherwise, $t_{k}^{(i)}=c_{k}^{(i)}$ if $\delta_{k}^{(i)}=0$. A survival model predicts the probability that some event $k$ occurs at time $T$ later than $t$, i.e., the so-called survival or event probability, which is denoted $S\left(t\right)=\text{Pr}\left(T>t\right)=1-\text{Pr}\left(t\leq T\right)$. In multi-event survival analysis, we want to predict the probability of different events occurring over time. For each patient $i$, this results in a matrix of survival functions, $\bm{S}^{(i)}=\{\bm{s}_{1}^{(i)},\bm{s}_{2}^{(i)},\ldots,\bm{s}_{K}^{(i)}\}$. Here, $\bm{s}_{k}^{(i)}$ represents the probabilities that patient $i$ does not experience event $k$ up to each time point $t$, i.e., $\bm{s}_{k}^{(i)}=(P(e_{k}^{(i)}>1),P(e_{k}^{(i)}>2),\ldots,P(e_{k}^{(i)}>T_{\text{max}}))$. See Fig. 1 for a visual overview.
@@
We follow Sechidis et al. [20] and first split the dataset into a training, validation and test set by 70%, 10%, and 20% using a stratified procedure. This ensures that the event times and censoring rates are consistent across the three sets and the $K$ events. The stratification procedure uses a random seed (0-9). After splitting the data, we impute missing values using the sample mean for real-valued covariates or the mode for categorical covariates based on the training set only. A list of missing values is available in the Supplement (S2 Table.). After imputing missing values, we encode categorical covariates using a one-hot encoding strategy and apply a $z$-score data normalization to speed up the training process and improve stability. No outlier detection was performed.
Our evaluation procedure consists of training five survival models on the training set and evaluating them on the holdout test set. For evaluation fairness, we do not perform hyperparameter optimization for specific models, but instead configure all models with sensible default parameters. A list of hyperparameters is available in the Supplement (S3 Table.). For deep learning models, we use a single hidden layer with 32 nodes. For methods that support it, we use early stopping during training to terminate the process if the validation loss does not improve for 10 consecutive epochs. All data preprocessing and experiments were implemented in Python 3.9. A list of software packages used is available in the source code repository.
@@
To evaluate both single-event and multi-event survival modeling approaches, we distinguish between models that predict the time to a single event and those that jointly model multiple event types. As single-event models, we implement the traditional Cox Proportional Hazards (CoxPH) model using a linear estimator [21], Random Survival Forests (RSF) [22], DeepSurv [23] and MTLR [24]. As a multi-event model, we implement the Multi-Event Network for Survival Analysis (MENSA) model [25]. To explain these models: CoxPH is a popular semiparametric method to fit a regression model to survival data. By adopting a multiplicative form for the contribution of several covariates to each individual’s event time, the CoxPH model is a powerful yet simple tool for assessing the simultaneous effect that different covariates have on the event times. CoxPH assumes proportional hazards, which means that the effect of covariates on the hazard function is constant over time. RSF extends decision trees for survival analysis and recursively splits training data based on some survival-specific criterion, e.g., log-rank splitting, with the goal of maximizing separation between event times between nodes. RSF can model nonlinear relationships between the covariates and the event, and does not assume proportional hazards. DeepSurv is a multilayer perceptron (MLP) based on the CoxPH model, where the risk score is a nonlinear function of the covariates. DeepSurv assumes proportional hazards. MTLR is a discrete survival model that estimates the individual survival function as a multinomial logistic regression model and does not assume proportional hazards. Lastly, MENSA is an MLP that learns the individual survival distribution for $K$ events as a mixture of Weibull distributions with a shared covariate layer to model dependencies between events. MENSA does not assume proportional hazards.
@@
To comprehensively assess the survival models, we select a range of evaluation metrics that capture three key aspects of model performance: discrimination, absolute/squared error, and calibration. Discrimination, usually measured by the concordance index (CI), is relevant when we want to identify which patients are likely to experience functional decline earlier versus later at the group level. Squared or absolute error metrics, like the mean absolute error (MAE), are used to assess how accurately the model predicts the time of functional impairment for individual patients. Finally, calibration measures how well the predicted probabilities match the observed frequencies of events at specific time points. For example, if a model predicts that a patient has an 80% chance of not experiencing significant functional decline before 100 days, then among 100 similar patients, about 80 should not experience significant functional decline by that time. We use the following metrics to assess discrimination, absolute/squared error, and calibration:
@@
CI: Harrell’s CI [26] measures discrimination performance by calculating the proportion of concordant pairs among all comparable pairs given predicted risk scores. A pair is considered comparable if the event order can be determined. Specifically, for a comparable pair $\{i,j\}$ with event times $e_{i}<e_{j}$ and $\delta_{i}=1$ (indicating that $i$ experienced the event), we calculate risk scores $\hat{\eta}_{i}$ and $\hat{\eta}_{j}$ as the negative median survival times. The median survival time is the time point at which the estimated survival probability drops to 0.5. If the predicted risk score for $j$ is greater than that for $i$ at the time $i$ experienced the event (while $j$ remains event-free), the pair is concordant. A CI of 0.5 indicates chance-level ranking. We also report Uno’s CI [27], which is a CI that uses a technique to correct for the bias introduced by censoring.
@@
BS/IBS: The Brier Score (BS) calculates the squared error between the predicted probability of the event and the Heaviside step function of the observed event. The integrated Brier score [28] (IBS) aggregates Brier scores over multiple time points to provide a single measure of the squared error.
@@
MAE: The MAE calculates the average absolute difference between the true event time and the predicted event time. Given an ISD, $S\left(t\mid\bm{x}_{i}\right)=\text{Pr}\left(T>t\mid\bm{x}_{i}\right)$, we calculate the predicted event time $\hat{t}_{i}$ as the median survival time [29]. Since our dataset contains censoring, we calculate the margin MAE (mMAE) as proposed by [13], which assigns a “best-guess” estimate to each censored patients using the Kaplan-Meier (KM) [30] estimator.
@@
D-calibration: Distribution calibration [13], also known as D-Cal, measures how well the predicted survival function, $\hat{S}(t)$, is calibrated for each event. D-calibration assesses this using a Pearson’s chi-squared test with $\alpha=0.05$. For any probability interval $a,b]\in[0,1]$, we define $D_{m}(a,b)$ as the group of patients in the dataset $D$ whose predicted probability of an event is in the interval $[a,b]$ [31]. A model is D-calibrated if the proportion of patients $|D_{m}(a,b)|/|D|$ is statistically similar to the amount bab-a. Intuitively, if a model is well-calibrated, the predicted probabilities match observed event frequencies – for instance, about 20% of patients predicted to have a 0.2 event probability should actually experience the event. In our experiments, we report the number of times each model is D-calibrated (at $p>0.05$) for each event across 10 experiments.
@@
We present our empirical results (Table 4 and Fig. 3) in terms of the aforementioned aspects with respect to model performance: discrimination, absolute/squared error, and calibration. Below, we discuss the results and reflect on their implications.